OpenStudy (anonymous):

Find nth term arithmetic sequence whose 1st term is 10 and whose 21st term is 0.

OpenStudy (anonymous):

solomon, any tips or hints?

OpenStudy (anonymous):

you still there by the way?

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

you got disconnected again?

OpenStudy (solomonzelman):

Finally.

OpenStudy (solomonzelman):

What do you know?

hartnn (hartnn):

use the formula \(\large a_n = a_1 + (n-1)d\)

OpenStudy (solomonzelman):

Oh my bad....

hartnn (hartnn):

a1 = 10 is given for 21st term put n= 21 and a21= 0

OpenStudy (anonymous):

so... I'm sorry, I'm kinda just learning this stuff myself because my teacher has alot of trouble speaking English :(

OpenStudy (anonymous):

solomonzelman, could you run that by me again?

OpenStudy (solomonzelman):

yes, I can show you an example or a couple.

OpenStudy (anonymous):

I'm kinda of a step-by-step kind of guy

hartnn (hartnn):

you noted down the formula ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so

OpenStudy (anonymous):

wat does a n, a1, n, and d stand for?

hartnn (hartnn):

ok, a_n is the n'th term a1 = 1st term d= common difference

hartnn (hartnn):

so, 21st term = 0 means we put n=21

hartnn (hartnn):

and an=0

OpenStudy (anonymous):

ok, so wait

OpenStudy (anonymous):

let me work this out

OpenStudy (anonymous):

so a1=10, right?

OpenStudy (solomonzelman):

yes

hartnn (hartnn):

yes, a1 = 1st term = 10 is given

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so wait

OpenStudy (anonymous):

I'm trying to think here

OpenStudy (anonymous):

if a1=10, wat would a n and n stand for?

hartnn (hartnn):

a_n is the n'th term like a1 = 1st term a2 = 2nd term a3 = 3rd term a4 = 4th term and so on

OpenStudy (solomonzelman):

\[a_n\]is the nth term, or in your case 21st.

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

so 0=10+(n-1)d?

OpenStudy (anonymous):

am I getting there yet?

OpenStudy (solomonzelman):

I think we can do an easier example to get it conceptually, real quick,

OpenStudy (anonymous):

ok

hartnn (hartnn):

0 = 10 + (21-1)d remember that n = 21, right ? because 21st term was given = 0

OpenStudy (solomonzelman):

Lets say you have an arithmetic sequence,\[1,4,7,11.....\]and you have to find the 10th term

OpenStudy (solomonzelman):

So far so good?

OpenStudy (anonymous):

35

OpenStudy (solomonzelman):

You have a formula, \[a_n=a_1+d(n-1)\]

OpenStudy (solomonzelman):

how, what did you do to get 35?

OpenStudy (anonymous):

I counted in my head

OpenStudy (solomonzelman):

Using formula or adding?

OpenStudy (anonymous):

adding

OpenStudy (solomonzelman):

Don't do this, becuase you won't be able to count 10000th term.

OpenStudy (anonymous):

though it generally only works for small units

OpenStudy (anonymous):

I know

OpenStudy (anonymous):

look

OpenStudy (anonymous):

a n=1+(10-1)d?

OpenStudy (anonymous):

is that any better?

OpenStudy (anonymous):

i don't about d though :(

OpenStudy (anonymous):

I don't know about d

OpenStudy (solomonzelman):

Don't do this, becuase you won't be able to count 10000th term.

OpenStudy (anonymous):

hmm?

OpenStudy (anonymous):

what do you mean, I am using the formula now

OpenStudy (solomonzelman):

I snapped again.

OpenStudy (anonymous):

I'm just trying to figure out know about the common difference

OpenStudy (anonymous):

and what it exactly stands for

OpenStudy (solomonzelman):